# Entanglement Dynamics in Monitored Systems and the Role of Quantum Jumps

### v2

@@ III.4 Models
III.5 Summary of Entanglement Dynamics in the No-Click Limit
@@
In the non-Hermitian Ising chain in a complex transverse field, Eq. (20), the entanglement entropy was found for weak monitoring $\gamma<\gamma^{\rm Ising}_{c}(h)$ to depend logarithmically on both time and subsystem size. The corresponding effective central charge $c_{\rm eff}$, obtained from the ansatz $S_{\rm nH}(\ell)=c_{\rm eff}\ln(\ell)$, could be obtained in closed form Turkeshi and Schiró (2023) and was shown to decrease as a function of both $\gamma$ and $h$ up to the critical line $\gamma^{\rm Ising}_{c}(h)$, above which the system was found to undergo an entanglement transition into an area-law phase, cf. Fig. 2(b). As for the SSH case, the entanglement transition for the Ising chains is directly related to a transition in the spectrum of non-Hermitian quasiparticles. The latter separates a critical gapless phase with vanishing imaginary part of the spectrum at a given point in the Brillouin zone, from a gapped phase Biella and Schiró (2021). In the next sections, we will discuss how the presence of quantum jumps modify the entanglement dynamics in these two models.
III.6 Relation to Prior Works
Here we position our work with respect to the current literature on monitored quantum systems. Non-interacting systems in particular have been discussed extensively, both numerically and analytically. Free fermions with global $U(1)$ symmetry, related to charge conservation at the single trajectory level, were shown to display a transition from a critical phase logarithmic scaling of entanglement entropy to an area-law for both quantum-jumps and quantum state diffusion (QSD) type of density-monitoring Alberton et al. (2021); Van Regemortel et al. (2021). A related transition was also found in the entanglement negativity in presence of boundary driving Turkeshi et al. (2022b). Our monitored SSH model falls a priori in the same class, since the global charge is conserved at the trajectory level, even though the monitored observable is not the total density which introduces some differences as we will discuss. This entanglement transition was first interpreted as a Kosterlitz-Thouless criticality based on a two-replica field theory and bosonization Buchhold et al. (2021). The robustness of the weak-monitoring sub-volume phase, and hence of the transition, was questioned in the case of projective measurements, where numerical evidence supported an area-law steady-state entanglement Coppola et al. (2022). This result has been theoretically understood within a replica field theory calculation Poboiko et al. (2023a) which identified connections with the physics of weak-localisation corrections in Anderson localisation. Later work has extended this analysis to higher-dimensions Poboiko et al. (2023b), finding a stable entanglement transition, as parallelly done within the QSD protocol, where non-trivial connections to multifractality were highlighted Chahine and Buchhold (2023).
When the symmetry is reduced to a discrete $Z_{2}$, corresponding to our Ising chain or Majorana fermions, a logarithmic to area law transition has been found both under QSD Turkeshi et al. (2021); Piccitto et al. (2022) and QJ Botzung et al. (2021); Turkeshi et al. (2022a); Paviglianiti and Silva (2023). In the latter case a phenomenological quasiparticle picture connected the latter to the no-click limit Turkeshi et al. (2022a). The replica field theory analysis, applied to a circuit model with temporal randomness and QSD monitoring also pointed towards a transition Jian et al. (2023), although with a different scaling of the entanglement Fava et al. (2023); Lóio et al. (2023).
With respect to those works we emphasize that the QJ protocol we consider here differs in one key aspect, namely noise distribution is state-dependent a priori and not prefixed, which makes the averaging over the noise, required to construct the replica field theory, non-trivial. Finally, for interacting systems QJ monitoring has been explored and shown to lead to a volume-to-area law transition Fuji and Ashida (2020).
@@ III.6 Relation to Prior Works
IV Entanglment Dynamics under Quantum Jumps
In this section we present our results for the QJ dynamics of the SSH and Ising chain model, focusing on the dynamics of the entanglement entropy. Unless specified otherwise, throughout this section we consider an initial product state of the form $|\Psi(0)\rangle=|0,1,0,1,\dots,0,1\rangle$ and open boundary conditions. We fix the hopping along the chains as units of energy and inverse time, $J=1$. We integrate the dynamics with a time-step $dt\propto 1/({4\gamma L})$. Throughout this work we consider a system of size $L=128$.
IV.1 Entanglement Dynamics in the Monitored SSH Model
We start discussing the entanglement dynamics for the monitored SSH. In Fig. 3 (a-c), we show the evolution of the entanglement entropy for a sample of quantum trajectories for increasing values of the monitoring rate $\gamma=0.5,1.5,2.5$, corresponding to the volume-law [panels (a,b)] and area-law [panel (c)] phase of the no-click limit, cf. Fig. 2. For comparison, we also plot the average entanglement entropy (obtained averaging over the different trajectories, according to Eq. (12)) as well as the value obtained in the non-Hermitian no-click limit. We see that, for weak monitoring, when the no-click dynamics displays a linear growth corresponding to the volume law phase, the stochastic dynamics due to quantum jumps suppress the entanglement very effectively and bring the system to a stationary state which is less entangled than in the no-click limit. In this regime, an agreement with pure non-Hermitian evolution is present only for short time scales. We also notice that for weak monitoring, trajectory-to-trajectory fluctuations are rather broad, and certain outliers follow the non-Hermitian evolution for longer times before ultimately deviating from it. On the other hand, in the regime of strong monitoring, when the non-Hermitian dynamics display an area law behavior, the role of quantum jumps is less drastic. While we still see strong fluctuations, the average entanglement entropy follows the no-click limit rather closely all the way to the stationary state. From these first results, we can already anticipate that QJs will have a disruptive effect on the volume law phase of the non-Hermitian SSH model. This can be understood by considering that the stochastic noise due to the jumps breaks the PT-symmetry of the non-Hermitian SSH model, which played a major role in protecting the volume law phase in the no-click limit Gal et al. (2023).
To better characterize the entanglement content in the jump dynamics, we consider the long-time steady state and study the scaling of the average entanglement entropy with the subsystem size $\ell$ as a function of the measurement rate $\gamma$. In Fig. 3 (d), we show that in the weak-monitoring regime, the volume law scaling of the no-click evolution is substituted by a much slower scaling compatible with a logarithmic growth, which eventually gives away for an area law phase upon increasing $\gamma$. Following the literature on monitored free fermions Alberton et al. (2021); Turkeshi et al. (2021, 2022a) we can use the phenomenological fit
$S^{\infty}_{\ell}=c_{\rm eff}\log\ell$ (21)
to obtain an estimate for the prefactor $c_{\rm eff}$, sometime called effective central charge $c_{\rm eff}$. As we show in the panels (e) and (f) of Fig. (3), this quantity vanishes as a function of both $\gamma$ and $h$, indicating a transition into an area-law phase.
[…]
As we see from panel (e), this transition occurs for values of $\gamma\simeq 1$, i.e., still within the volume law phase of the no-click problem (See Fig. 2). Furthermore, the critical point $\gamma_{c}$ decreases weakly for $h<2$, as opposed to the no-click result. On the other hand, we note that the effective central charge is non-monotonous as a function of $h$, vanishing around $h\simeq 2$. This result can be understood from the fact that the quasiparticle band of the non-Hermitian Hamiltonian becomes flat around that value Gal et al. (2023) and, as a consequence, the QJ dynamics at short-intermediate times slows down. We emphasize that the entanglement transition between log and area law scaling discussed above could well turn into a crossover upon increasing system sizes through a weak-localization mechanism similar to what was recently discussed in the case of free fermions with $U(1)$ symmetry and projective measurements Poboiko et al. (2023a). Although affected by the inevitable finite-size effects, our results show that the volume law phase of the non-Hermitian SSH is not stable when including quantum jumps. As compared with the phase diagram of the no-click limit, we see that not only does the volume-law turn into a logarithmic phase but also that the emerging area law phase is significantly more extended throughout the phase diagram, which is therefore qualitatively and quantitatively affected by the inclusion of quantum jumps.
IV.2 Entanglement Dynamics in the Monitored Ising Chain
We now move to the monitored Ising chain and present our results for the entanglement entropy dynamics under QJs. These complete previous results obtained in Ref. Turkeshi et al. (2022a); Paviglianiti and Silva (2023). As done before for the SSH case, we start in Fig. 4 presenting the dynamics of the entanglement entropy for a certain number of stochastic trajectories and compare to the average value and the value obtained in the no-click limit.
As a first observation, we note that in the weak monitoring phase, $\gamma=0.5,\ h=0.4$ [panel (a)], the non-Hermitian result follows quite closely the average entanglement entropy and the fluctuations appear to remain modest in size. At long times, we see a deviation, namely, the non-Hermitian dynamics reach a maximum and then decrease, while the QJs attain a stationary state value, which still contains fluctuations. As the value of the monitoring rate $\gamma$ increases [cf. panel (b) for $\gamma=1.5,\ h=0.4$], we observe a similar trend, namely that quite generically, the agreement between no-click and QJs is good at short and intermediate times, while at long-times quantum jumps can stabilize a more entangled state than in the pure no-click non-Hermitian case.
After this qualitative discussion, we present in Fig. 4 (d-f) the scaling of the steady-state entanglement entropy as a function of the subsystem size $\ell$, for different values of the parameters. In panel (d) we confirm that the weak monitoring phase has a logarithmic dependence of the entanglement entropy, which evolves into an area law upon increasing $\gamma$. Following previous works we extract an effective central charge $c_{\rm eff}$ by fitting a logarithmic scaling of the entanglement entropy in the stead state, cf. Eq. (21). In Fig. 4 (e,f) we plot this quantity as a function of $\gamma$ and $h$. We see that increasing $\gamma$ at fixed $h$ drives a transition into an area law phase when $c_{\rm eff}=0$, in qualitative analogy to the no-click limit. The critical point $\gamma_{c}(h)$ decreases weakly with the transverse field [see panel (e)], a result which is qualitatively similar to the no-click case where however this effect is more pronounced, cf. Fig. 2. If we instead fix the value of the monitoring rate and increase the transverse field $h$ we see in panel (f) a rather different behavior: for small $\gamma$ the effective central charge has a maximum and then decreases with $h$, while for larger values of $\gamma$ we still see an effect similar to the no-click limit, with $c_{\rm eff}$ vanishing at a critical $h$. This has an interesting consequence, already noticed in Ref. Paviglianiti and Silva (2023), namely that for large $h>1$ the monitored problem shows a sub-volume logarithmic scaling of the entanglement entropy in a region where according to the no-click evolution the system should be in the area law. Here we find similar results as we show in Fig. 4 (e,f): both at large $h>1$ and at large $\gamma>2.0$ the system still displays a finite $c_{\rm eff}>0$, while the no-click limit would predict an area law phase.
The outcome of this comparison confirms that, quite generally, the QJ dynamics in the Ising chain retain many features of the non-Hermitian phase diagram. However, the phase transition into the area law is renormalized concerning the no-click limit value. In particular, quantum jumps enhance the entanglement entropy to a logarithmic law in regions of the phase diagram where the no-click dynamics would predict an area law. These results highlight a compelling difference between the SSH and Ising models. While the former QJs have severe effects on the weakly monitored phase, deteriorating the volume-law phase to a logarithmic phase, the latter QJs show a marked contribution in the area-law phase of the non-Hermitian Hamiltonian, cf. for a measurement-altered analysis Paviglianiti et al. (2023); Murciano et al. (2023). We will return to this point in Sec. VII.
To summarize, in this Section we have shown that QJs affect the dynamics of entanglement entropy in rather different ways for our two models. Similar conclusions can be drawn by looking at other non-linear functions of the state, e.g., the statistics of the monitored density, that we discuss in Appendix A. In the following sections, we will provide a theoretical framework to understand this difference and the deviations with the no-click limit.
V Statistics of Quantum Jumps
The results of the previous Section have shown that QJs act as a strong perturbation concerning the dynamics of the non-Hermitian SSH, while they appear less relevant for the Ising chain. In particular, we have shown that the non-Hermitian dynamics leave their fingerprints on the stochastic evolution in the Ising case, while it appears to be washed away for the SSH model. In this Section, we try to understand whether such a difference arises from different statistical properties of QJs in those two models. Indeed, one possible scenario to explain the different roles of quantum jumps could be that the statistics of quantum jumps become particularly broad and, as a consequence, jumps are rare and diluted, leaving more space for the non-Hermitian evolution to control the physics. As we will see in this Section, this naive expectation is incorrect.
To study how QJ are distributed in time, we compute their Waiting-Time Distribution (WTD). This quantity naturally emerges in the numerical implementation of stochastic QJ dynamics beyond the first-order Monte Carlo schemes Daley (2014). Furthermore, WTDs have been studied in the early days of quantum jumps, motivated by resonance fluorescence spectra Cohen-Tannoudji and Dalibard (1986); Carmichael et al. (1989), and have found a multitude of applications from solid-state quantum information Delteil et al. (2014) to quantum transport Brandes (2008); Albert et al. (2012); Landi (2021, 2023); Coppola et al. (2023); Landi et al. (2023) to laser cooling Bardou et al. (1994).
Figure 5: Waiting Time Distribution for the SSH model obtained numerically for different values of $\gamma$. In the left panel (a) the entire dynamics is considered whereas in the right panel (b) only QJs occurring in the long-time steady state are sampled. We see that in the latter case a clear Poisson law emerges, while short-time jumps are affected by the initial condition and display a longer average waiting time.
We start by discussing the SSH case for which we plot the WTD (obtained numerically) in Fig. 5. We note from panel (a) that the WTD displays an exponential decay typical of a Poisson law at weak monitoring. In contrast, upon increasing the monitoring rate, we see the development of nonexponential tails, giving rise to longer waiting times. This fact can partially explain the agreement between QJs and non-Hermitian dynamics at short times, see for example in Fig. 3. While the results in panel (a) have been obtained by sampling the jumps over full-trajectories, in panel (b), we consider only the jumps occurring at long times, i.e., in the steady state, for which we see a clear Poisson law emerging for all monitoring rates. We can better understand the origin of the tails in the short-time WTD by recalling that this distribution is controlled by the decay of the norm $\mathcal{N}(t)\equiv||e^{-iH_{\mathrm{eff}}t}|\Psi\rangle||$, cf. Eq. (8), which reads in general $\mathcal{N}(t)=\exp(-K(t)t)$. In the case of the SSH model the observable $K(t)$ driving the decay of the norm is $K_{\mathrm{SSH}}(t)=2\gamma\sum_{i}[n_{A,i}(t)+1-n_{B,i}(t)]$, where $n_{A/B,i}(t)=\langle\Psi(t)|c_{A/B,i}^{\dagger}c_{A/B,i}|\Psi(t)\rangle$. At short times, the observable $K_{\mathrm{SSH}}(t)$ displays significant temporal fluctuations due to the chosen initial condition, and the WTD shows longer tails. On the other hand, at long times one obtains a Poisson law
$P_{WT}(\tau;t\rightarrow\infty)\sim e^{-K_{\infty}\tau}$ (22)
whose average waiting time is scaling as $1/\gamma L$ and thus increases as $\gamma$ decreases, cf. Fig. 6 (b). A Poisson distribution of QJ is typical for systems in which the backaction term is a constant of motion. In the SSH setup, this is not the case, and in fact, we note that this WTD is a property of the steady-state jump dynamics due to the observable $K(t\rightarrow\infty)$.
Similarly, we compute the WTD for the Ising chain ; see Fig. 6 (a). Even in this case, we find that jumps are distributed according to a Poisson law. To understand this result, it is important to recall that jumps try to refill empty sites and keep the total number of particles effectively constant. Thus, $K(t)$ is centered around the average number of particles with small fluctuations. Interestingly, both models feature the same average waiting time of QJs, as shown in Fig. 6 (b), and given by $\tau\sim 1/\gamma L$.
Figure 6: Waiting Time Distribution for the monitored Ising chain obtained numerically for different values of $\gamma$, showing a perfect agreement with a Poisson law. In the bottom panel, we compare the average waiting time of quantum jumps in the two models, showing perfect agreement. In both cases $\overline{\tau}\sim 1/\gamma L$.
Figure 7: Average change to the entanglement entropy due to non-Hermitian evolution (panel a) and QJs (panel b). These averages are obtained from the joint distributions, $P(\Delta S_{\rm qj},S),\ P(\Delta S_{\rm nh},S)$, thus are a function of the entanglement entropy $S$ and we represent in these plots the dependence from subsystem size $\ell$.
VI Statistics of Entanglement Gain and Loss
Figure 8: Stochastic model for entanglement dynamics — Panel (a), (b) total rate of entanglement entropy growth obtained, for the monitored Ising chain, from the average entanglement gain and loss, according to Eq. (24). We see the rate $\Phi(S,\ell)$ vanishes at a value $S_{\infty}(\ell)$ which identifies the steady-state entanglement (circled points). Panel (c) compares $S_{\infty}$ obtained from the stochastic model with the long-time limit of the entanglement entropy obtained from the full QJ dynamics, showing a perfect agreement.
The study of the statistics of waiting times for quantum jumps discussed in the previous Section reveals a rather common pattern for the two models, namely, at long times, the WTD of the jumps flows towards a universal Poisson law, independent of the state, with an average waiting time $\bar{\tau}\sim 1/\gamma L$. This suggests that it is not the time distribution or the number of jumps that control their impact on the entanglement structure of the state. To better understand the intertwined role of QJs and non-Hermitian dynamics, we introduce a new metric that looks at the statistics of the change in entanglement entropy after and in-between quantum jumps. Specifically, we sample throughout a quantum trajectory (see Fig. 1 (b)) the changes in entanglement entropy at each jump event and in between the jumps (due to the non-Hermitian evolution) and repeat this sampling for different trajectories to collect two histograms $P(\Delta S_{\rm qj})$ and $P(\delta S_{\rm nh})$. We have plotted these histograms in Fig. 1 (c-d) for the monitored Ising chain. The first interesting observation is that the QJ distribution is strongly peaked around $\Delta S_{\rm qj}=0$ and has broad asymmetric tails, suggesting that the vast majority of quantum jumps do not substantially change the entanglement entropy. In contrast, rare quantum jumps are responsible for more significant changes, and this can be also observed at the level of the trajectory in Fig. 1 (b). The asymmetry of the distribution is also interesting to note since it means that a single QJ is more likely to reduce the entanglement entropy, even though jumps that increase it are also possible. For what concerns the entanglement changes during the non-Hermitian evolution in between quantum jumps, we see that the distribution appears centered around a slightly positive value and with a slight asymmetry in the tails, suggesting that the non-Hermitian evolution is primarily responsible for the growth of the entanglement.
In the same plots in Fig. 1 (c-d) we also consider the joint probability of observing a change $\Delta S_{\rm qj}$ (or $\delta S_{\rm nH}$) in correspondence of a particular value of entanglement entropy $S$, i.e., $P(\Delta S_{\rm qj},S),\ P(\delta S_{\rm nH},S)$ which are obtained by binning stochastic events (QJs or non-Hermitian evolution) according to the entanglement content of the state they act on. The key idea here is to understand whether QJs impact differently states which are highly or weakly entangled. As we see in Fig. 1 (c-d), the tails of the histogram for the QJs broaden up (at least for $\Delta S_{\rm qj}<0$), indicating that highly entangled states are indeed more fragile and prone to be affected by rare QJs.
To condense the rich information contained in the full statistics of entanglement changes we now focus on the first moment of those distributions, corresponding to the average change to the entanglement entropy due to quantum jumps $\overline{\Delta S}_{\rm qj}(S,\ell)$ and to the non-Hermitian evolution $\overline{\delta S}_{\rm nH}(S,\ell)$. We emphasize that these averages  [^1] With a slight abuse of notation, we indicate the average over the gain/loss statistics with an overbar, as for the average over trajectories. depend on the average entanglement entropy $S$ since they are evaluated over the joint distributions $P(\Delta S_{\rm qj},S)$ and $P(\delta S_{\rm nH},S)$, i.e.
$\overline{\Delta S}_{\rm qj}(S,\ell)\equiv\int_{\Delta S_{\rm qj}}\,\Delta S_{\rm qj}\,P(\Delta S_{\rm qj},S)$ (23)
and similarly for $\overline{\delta S}_{\rm nH}(S,\ell)$. In Fig. 7 (a-b), we plot these first moments for the monitored Ising chain (at a particular point of the phase diagram) for different subsystem sizes $\ell$. In panel (a), we see that the change due to non-Hermitian evolution is on average positive, $\overline{\delta S}_{\rm nH}>0$, i.e., it induces an entanglement entropy gain. This gain is substantial for weakly entangled states, then decreases slowly with the entanglement content. In panel (b), we plot the change due to QJs, normalized with respect to the average waiting time $\overline{\tau}$. As shown in Appendix B, this ratio remains well defined upon increasing the system size $L$, despite the waiting time vanishing as $\overline{\tau}\sim 1/L$. This is consistent with the idea that a single quantum jump can affect the entanglement entropy of a quantity $O(1/L)$. From Fig. 7(b), we see that $\overline{\delta S}_{\rm qj}/\overline{\tau}$ is negative and decreases with $S$, i.e. QJs on average induce an entanglement entropy loss. They do so the more the state onto which they act is entangled. Interestingly, both averages develop a subsystem-size $\ell$ dependence above a certain threshold entanglement entropy value. A natural question that we adress next is how these average entanglement gains and losses are connected to the entanglement dynamics under QJs and its phase transition.
VI.1 Classical Stochastic Model for Entanglement Dynamics
We now present a classical stochastic model that builds upon the entanglement gain and loss statistics discussed in the previous section and gives a physical picture of the entanglement dynamics under QJs. In particular, we model the entanglement evolution as a random walk with random drift and partial resetting Evans and Majumdar (2011); Evans et al. (2020); Turkeshi et al. (2022a), which is sketched in Fig. 1 (e) and that we will now describe. We use the previous analysis showing that the waiting time $\tau$ between jumps follows a Poisson law, which means that the random variable can be expressed as $\tau=-\log(r)/(K_{\infty}(\gamma))$ were $r\in[0,1]$ is a random number drew for each jump. Then, during this time $\tau$, we model the change of entanglement due to the non-Hermitian evolution by $\Delta S_{\rm nH}=\tau\leavevmode\nobreak\ \delta S_{\rm nH}$, where $\delta S_{\rm nH}$ is picked with a probability $P(\delta S_{\rm nH},S)$ from the previously computed joint distribution of the effective non-Hermitian slopes. After the time $\tau$, the instantaneous jump happens, and we model the change it has on the entanglement by $\Delta S_{\rm qj}$, which is drawn with a probability $P(\Delta S_{\rm qj},S)$ from the corresponding probability distribution. In both cases, the entropy $S$ corresponds to the entanglement before the event (i.e., before the jump or before the non-Hermitian evolution). The probability distributions we use here are taken from the exact sampling of quantum trajectories, binned according to the entanglement entropy content.
[…]
$\frac{dS}{dt}=-\overline{\delta S_{\rm nH}}(S,\ell)-\overline{\Delta S_{\rm qj}}(S,\ell)/\overline{\tau}\equiv\Phi(S,\ell)$ (24)
This is a straightforward rate equation for the dynamics of the average entanglement entropy $S$, which appears to be controlled by a balance between average entanglement gain $\overline{\delta S_{\rm nH}}(S,\ell)$ and loss $\overline{\Delta S_{\rm qj}}(S,\ell)$. The fact that these gain/loss rates depend on the value of the entanglement entropy itself, as previously discussed, is crucial here and results in a non-trivial flow of the entanglement entropy with time, encapsulated in the function $\Phi(S,\ell)$ defined in Eq. (24). In Fig. 8 (a-b), we plot $\Phi(S,\ell)$ for the Ising chain (at two particular points of the phase diagram). We see that $\Phi(S,\ell)$ vanishes at a value $S=S_{\infty}(\ell)$, which by construction identifies the steady-state entanglement and which is obtained when the gain due to non-Hermitian evolution and the loss due to QJs perfectly balance each other, i.e.
This fixed point is attractive, i.e., depending on the initial condition, it is approached either from the low-entangled branch where $\Phi(S,\ell)>0$ or from the high-entangled one where $\Phi(S,\ell)<0$.
We can now validate this stochastic model by extracting the steady-state entanglement $S_{\infty}$ and its subsystem size $\ell$ dependence from Eq. (25) for the monitored Ising chain and compare it with the full QJ results discussed in Sec. IV. As shown in Fig. 8(c), the agreement is really good both at weak monitoring, where it reproduces the logarithmic scaling of the entanglement entropy, and in the area law. More importantly, the deal is not limited to the steady-state, but it extends to the full dynamics, as we have shown in Fig. (1) (f). A similar agreement for the steady-state and dynamics of entanglement entropy is also obtained for the monitored SSH model (Appendix D).
We emphasize that the classical stochastic model leading to Eq. (24) is built upon the entanglement gain and loss statistics obtained from the full QJ dynamics. Nevertheless, it still represents a considerable simplification with respect to the total quantum stochastic dynamics, and its agreement points towards the universality of the effect of quantum jump and non-Hermitian evolution on entanglement, which is intriguing and worth exploring further. More importantly, as we will discuss now, it provides a new angle to analyze the dynamics of entanglement under QJs.
VII Steady-State Entanglement Balance: Non-Hermitian vs Jumps Contributions
In this Section, we use the insights from the classical stochastic model to look at the results of the full QJ dynamics presented in Sec. IV from a different perspective. Indeed, we have learned that the steady-state entanglement entropy under QJ monitoring arises from the competition between gain due to the non-Hermitian dynamics and losses due to the jumps. Here, we will discuss the relative balance of these two contributions across the phase diagrams of our two models. This will allow us to shed light on the origin of the entanglement scaling in the two models and the mutual role of jumps and non-Hermitian dynamics. To qualify for a meaningful comparison, we choose values of $\gamma$ to give the same average waiting time and thus the same number of jumps per trajectory in average.
Figure 9: Steady-State Entanglement Balance for the Monitored SSH chain: Average entanglement gain $\overline{\delta S_{\rm nH}}$ and loss $-\overline{\delta S_{\rm qj}}/\overline{\tau}$, as a function of the average entanglement entropy $S$ and different subsystem sizes $\ell$ (error bars are smaller than the size of the dots). In the weak monitoring phase, Panel (a), the contribution due to QJs scale logarithmically with $\ell$(see Inset), while the term arising from the non-Hermitian evolution is practically flat. In the strong monitoring phase, corresponding to the area law, both contributions depend on $\ell$. The crossing point corresponds again to the zero of $\Phi(S,\ell)$ i.e., to the steady state entanglement $S_{\infty}$.
VII.1 Monitored SSH
Let us start discussing the SSH model at weak monitoring $\gamma=0.2,h=1.0$, corresponding to the log-phase discussed in Sec. IV. In Fig 9 (a) we plot separately the average entanglement gain $\overline{\delta S_{\rm nH}}$ and loss $-\overline{\Delta S_{\rm qj}}$ (with a minus sign since the steady-state condition from Eq. (25) now reads $\overline{\delta S_{\rm nH}}=-\overline{\Delta S_{\rm qj}}/\overline{\tau}$, leading to a crossing between the two rates). We see clearly that the jump contribution depends strongly on subsystem size ; the bigger the subsystem, the less QJs decrease the entanglement on average. Interestingly, the subsystem size dependence of this contribution looks logarithmic, as shown in the inset of panel (a). On the other hand, the contribution of the non-Hermitian evolution depends weakly on both $\ell$ and $S$, at least for the values of entanglement entropy, which are allowed to be explored by the jumps, i.e., around the crossing point which corresponds to the steady-state entanglement $S_{\infty}$.

### v3

@@ Acknowledgements.
Appendix A Summary of Entanglement Dynamics in the No-Click Limit
@@
In the non-Hermitian Ising chain in a complex transverse field, Eq. (11), the entanglement entropy was found for weak monitoring $\gamma<\gamma^{\rm Ising}_{c}(h)$ to depend logarithmically on both time and subsystem size. The corresponding effective central charge $c_{\rm eff}$, obtained from the ansatz $S_{\rm nH}(\ell)=c_{\rm eff}\ln(\ell)$, could be obtained in closed form Turkeshi and Schiró (2023) and was shown to decrease as a function of both $\gamma$ and $h$ up to the critical line $\gamma^{\rm Ising}_{c}(h)$, above which the system was found to undergo an entanglement transition into an area-law phase, cf. Fig. 8(b). As for the SSH case, the entanglement transition for the Ising chains is directly related to a transition in the spectrum of non-Hermitian quasiparticles. The latter separates a critical gapless phase with vanishing imaginary part of the spectrum at a given point in the Brillouin zone, from a gapped phase Biella and Schiró (2021).
Appendix B Statistics of Jumps Waiting Times
In this Appendix we discuss how QJs are distributed in time, i.e. we compute their Waiting-Time Distribution (WTD). This quantity has been studied in the early days of quantum jumps, motivated by resonance fluorescence spectra Cohen-Tannoudji and Dalibard (1986); Carmichael et al. (1989), and have found a multitude of applications from solid-state quantum information Delteil et al. (2014) to quantum transport Brandes (2008); Albert et al. (2012); Landi (2021, 2023); Coppola et al. (2023); Landi et al. (2023) to laser cooling Bardou et al. (1994) and is also a key quantity in this work.
In general, computing the waiting time distribution is a non-trivial task Landi et al. (2023). There is, however, a special case where this can be done straightforwardly. Whenever the imaginary part of the non-Hermitian Hamiltonian commutes with the real part (and that the initial state is an eigenstate of the non-Hermitian part), then one can conclude that the norm decay is exponential and therefore so the cumulative waiting time-distribution, $F=1-\exp\left(-K\tau\right)$ where $K=\sum_{i}\langle L^{\dagger}_{i}L_{i}\rangle$. In this case, the time for the next jump can be directly obtained as a function of a random number $r$, as $\tau=-\frac{1}{K}\mbox{ln}r$. Such an instance occurs for a system with local density monitoring and conservation of global particle number Alberton et al. (2021); Coppola et al. (2022).
Here this quantity has to be extracted numerically, but emerges naturally from the numerical implementation of stochastic QJ dynamics beyond the first-order Monte Carlo schemes Daley (2014). We start recalling that the WTD is obtained from the decay of the norm $\mathcal{N}(t)\equiv||e^{-iH_{\mathrm{eff}}t}|\Psi\rangle||$, which reads in general $\mathcal{N}(t)=\exp(-K(t)t)$ where $K(t)$ represents the back-action associated to the non-Hermitian Hamiltonian $H_{\mathrm{eff}}$, i.e. $K(t)=\sum_{i}\langle L^{\dagger}_{i}L_{i}\rangle$. At short times the rate of decay of the norm can display fluctuations (depending on the model and the initial state) which can result in biasing the WTD. However at sufficiently long-times one expect a Poisson law
$P_{WT}(\tau;t\rightarrow\infty)\sim e^{-K_{\gamma}^{\infty}\tau}$ (22)
at least for systems where the monitoring process is local on each sites (where $K_{\gamma}^{\infty}$ denotes the steady state value of $K(t)$ which eventually depends on the monitoring rate $\gamma$).
Our numerical analysis on the monitored Ising and SSH models confirm the expectation that QJs are Poisson distributed. The SSH model displays more pronounced tails in the short-time WTD, caused by fluctuations due to the initial state. Indeed for the SSH model the observable $K(t)$ driving the decay of the norm is $K_{\mathrm{SSH}}(t)=2\gamma\sum_{i}[n_{A,i}(t)+1-n_{B,i}(t)]$, where $n_{A/B,i}(t)=\langle\Psi(t)|c_{A/B,i}^{\dagger}c_{A/B,i}|\Psi(t)\rangle$. At short times, the observable $K_{\mathrm{SSH}}(t)$ displays significant temporal fluctuations due to the chosen initial condition, and the WTD shows longer tails. On the other hand in the Ising model jumps try to refill empty sites and keep the total number of particles effectively constant. Thus, $K(t)$ is centered around the average number of particles with small fluctuations. A Poisson distribution of QJ is typical for systems in which the backaction term is a constant of motion. Interestingly, both models feature the same average waiting time of QJs, as shown in Fig. 9 (b), and given by $\tau\sim 1/\gamma L$.
Figure 9: Average waiting time of quantum jumps in the Ising and SSH models, showing perfect agreement. In both cases $\overline{\tau}\sim 1/\gamma L$.
Appendix C System size scaling of entanglement loss
In the main text we introduced a classical stochastic model for the entanglement entropy dynamics, which builds upon the entanglement gain and loss statistics. In particular, the role of QJs is to induce an entanglement loss that, in the stationary state, is balanced by the gain provided by the non-Hermitian evolution, as described in Eq. (20) of the main text. This equation divides the entanglement loss by the average waiting time $\overline{\tau}$, which considers the instantaneous nature of QJs. As we have seen in Appendix. B, the average waiting time scales as $1/L$, which raises the question of the stability of the steady-state condition in the large system size limit. In this Appendix, we provide evidence supporting the statement that the average entanglement loss due to the jump scales as $1/L$, so the ratio with the averaging waiting time remains finite when $L\rightarrow\infty$. In Fig. 10, we plot, for the monitored Ising chain at a representative value of the parameters, the average entanglement loss divided by the waiting time, $\overline{\Delta S_{\rm qj}}(S,\ell)/\overline{\tau}$ as a function of $S,\ell$ and different system sizes. We see that the data corresponding to different system sizes collapses onto each other for each subsystem sizes $\ell$, which demonstrate this invariance of the ratio $\overline{\Delta S_{\mathrm{qj}}}/\overline{\tau}$.
@@ Appendix C System size scaling of entanglement loss
Appendix D Quantum jumps from highly entangled initial states
Figure 11: Analysis of trajectories obtained from a highly entangled initial state in the SSH model. Panel (a) : Entanglement Dynamics in the monitored SSH chain under QJ dynamics with an initial state highly entangled (namely a steady state of the no-click limit). In the inset we show the conditioned version of the distribution $\Delta S_{\rm qj}$ with respect to the entanglement content when jumps happen. Panel (b) : First moment of the conditioned distribution $\delta S_{\rm nH}$ in function of the entanglement content $S$. This quantity is obtained from these trajectories starting from the highly entangled initial state and which allow to probe the large entanglement behaviour of the distribution
In this Appendix we provide evidence to support the conjecture that highly entangled gaussian states are more fragile to measurements than weakly entangled ones. To this extent, we consider the monitored SSH chain starting from an initial condition corresponding to the long-time limit of the associated non-Hermitian Hamiltonian, which is known to support volume-law entangled states for small values of $\gamma$ Gal et al. (2023). In Fig. 11 (a) we show a sample of trajectories for the entanglement entropy, all converging towards a steady-state value with low entanglement (and equal to the steady-state reached from a product state initial condition). In other words, the system under monitoring cannot sustain volume-law entanglement. In panel (b) we plot the conditional distribution of entanglement loss $P(\Delta S_{\rm qj},S)$, which now displays a broadening of its tails indicating that the role of jumps become more relevant. In particular, we see that the probability of a large entanglement loss due to jumps at atypical (i.e. high) entanglement content increases, which explains why the initial state entanglement cannot be preserved. In panels (c) we present the same analysis given in the main text on the first moment of the gain conditional distribution $\overline{\delta S_{\rm nH}}$, but for this particular initial condition which allows us to probe entanglement content of higher values. We note that in the region of the steady state (at relatively low entanglement $S\leq 2$ delimited by the dashed line), the average entanglement gain from non-Hermitian dynamics is essentially independent from subsystem size $\ell$, a behavior similar to what observed in Fig. LABEL:fig:entanglement_breakdown_ssh for a different initial condition. However, as the system dynamics explores larger values of entanglement entropy ($S\geq 2$) we see important subsystem size effects in $\overline{\delta S_{\rm nH}}$. This behaviour is usually hidden when starting from a low entangled initial condition because the jumps are confining the dynamics in this lowly entangled space.
Appendix E Monitored Ising Chain at Large Field